### v1

@@ 5.2.2 Reinforcement Learning with OT Supervision
$r_{t}=r_{OT}(A_{t},P^{*}_{t})+\cdot r_{smooth}(A_{t},A_{t-1})+\cdot r_{cost}(A_{t},s_{t})$ (3)
@@
The smoothness term $r_{smooth}(A_{t},A_{t-1})=-\|A_{t}-A_{t-1}\|_{F}^{2}$ penalizes large changes between consecutive allocation matrices, promoting temporal continuity and preventing abrupt load redistributions that can cause system instability. This term is crucial for ensuring that regional servers experience gradual load transitions rather than sudden spikes or drops.
The cost term $r_{\text{cost}}(Q_{t})=-\frac{\|Q_{t}\|_{1}}{Q_{\text{max}}}$ utilizes the aggregate backlog across regional servers as a real-time proxy for system responsiveness, where $Q_{t}$ represents the vector of pending task counts and $Q_{\text{max}}$ denotes the maximum tolerable queue capacity. By continuously penalizing queue accumulation proportional to the normalized L1 norm, this reward component incentivizes prompt workload distribution before server congestion develops.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{PPO}}+\mathcal{L}_{\epsilon}(s,\epsilon)+\mathcal{L}_{s}(s,\epsilon)$ (5)
@@
This design ensures that TORTA operates within the theoretical trade-off region where reduced switching costs outweigh single-timeslot optimality losses. The constraint parameters $\epsilon$ and $s$ must satisfy a specific relationship to guarantee theoretical performance advantages over any reactive method: $\frac{1-1/s}{\epsilon}>\frac{L_{R}+\beta L_{P}}{K_{0}}$, where $L_{R}$, $L_{P}$ are Lipschitz constants of response time and power cost functions, and $K_{0}$ is the baseline switching cost of reactive methods. The rigorous mathematical derivation of this relationship, detailed training procedures, and the formal proof are provided in Appendix A and  B.
5.3 Micro-Level Allocation
@@
5.3.1 Dynamic Server Activation Strategy
@@
5.3.2 Greedy Task-Server Matching
@@
where $Recent_{tasks}^{(s)}$ represents tasks recently processed by server $s$, and $Similarity(\cdot,\cdot)$ measures task similarity based on model type, input characteristics, and processing requirements.
@@
/* Phase 1: Macro-Level Regional Allocation */
1
/* Demand and Supply Normalization */
@@ 5 $Prob_{i\rightarrow j}\leftarrow$ Row-normalize $P^{*}$;
6
/* Reinforcement Learning Optimization */
7 $F_{t}\leftarrow$ Predictor$(U_{t-K:t},Q_{t-K:t},H_{t-K:t})$;
@@ 8 $A_{t}\leftarrow\pi_{\theta}(s_{t},F_{t},A_{t-1})$;
/* Phase 2: Micro-Level Server Selection */
@@
       /* Dynamic Server Activation */
11       $N_{target}^{(r,t)}\leftarrow$ Activate target servers;
       /* Task-Server Matching */
12       $\mathcal{T}_{r}\leftarrow$ Get and sort regional tasks by urgency;
13
@@ 12       $\mathcal{T}_{r}\leftarrow$ Get and sort regional tasks by urgency;
15             $Score(i,s)\leftarrow$ Compute scores for all servers $s$;
@@ 16             $s^{*}\leftarrow\arg\max_{s}Score(i,s)$;
17             $X_{i,r,s^{*}}^{t}\leftarrow 1$, $Util_{s^{*}}\leftarrow Util_{s^{*}}+ResourceReq_{i}$;
18
19
/* System State Update */
20 Update $U_{t+1}$, $Q_{t+1}$, $A_{t-1}\leftarrow A_{t}$;
21
return $X_{t}$, $s_{t+1}$
@@ 6.1 Experimental Setup
Network Topology and Infrastructure. We evaluate our scheme across four distinct network topologies that represent different scales and geographical distributions of GPU infrastructure. The topologies range from small-scale regional deployments to large-scale global distributions (Table 1.a), enabling us to assess scalability and performance across various infrastructure configurations.
Each server cluster contains heterogeneous GPU resources with different capabilities and capacities (see Table 1.b). The infrastructure setup is designed to reflect real-world deployment scenarios where different GPU types are optimized for specific workload characteristics.
@@ 6.2 Evaluation Metrics
Table 1: Experimental Infrastructure Configuration
@@
Topo. Node Lat.(ms)
Abilene 12 25
Polska 12 45
Gabriel 25 80
Cost2 32 150
@@
6.3 Experimental Analysis
6.3.1 Response Time Performance Analysis
The response time distributions shown in Figure 8 demonstrate the effectiveness of our TORTA sheme across four different network topologies. The results indicate that TORTA consistently achieves the fastest average task completion times across all tested configurations. Specifically, TORTA obtains mean response times of 16.39s, 19.31s, 17.58s, and 19.19s for Abilene, Cost2, Gabriel, and Polska topologies respectively, representing significant improvements over baseline methods.
@@ 6.3.1 Response Time Performance Analysis
Figure 9: Power and switching cost comparison across different network topologies. Power costs are measured in thousands of dollars (\$K), while switching costs represent the relative overhead (normalized units).
@@
6.3.2 Cost Optimization and Economic Efficiency
Figure 9 presents a comprehensive cost analysis comparing TORTA with baseline methods across four network topologies. The analysis encompasses two critical cost components: power consumption costs and task switching overhead. TORTA consistently achieves the lowest power costs across all topologies, with reductions of approximately 10-15% compared to baseline methods. This improvement stems from the optimal transport component in our macro-level allocation, which explicitly considers regional electricity pricing differences when distributing tasks across geographical regions.
The power cost optimization is particularly evident in the Cost2 and Gabriel topologies, where TORTA achieves power costs of 11.1\$K and 10.7\$K respectively, compared to significantly higher costs for baseline methods. The optimal transport formulation enables TORTA to leverage geographical cost variations by preferentially routing computationally intensive tasks to regions with lower electricity prices, while maintaining acceptable network latency constraints. This cost-aware allocation strategy represents a key advantage of incorporating economic factors directly into the optimization objective.
Regarding switching costs, TORTA demonstrates substantial reductions across all topologies, with particularly notable improvements in Abilene and Cost2 configurations. The reinforcement learning component explicitly models state transition costs in its reward function, encouraging allocation decisions that minimize expensive server state changes and task migrations. Additionally, the micro-level server selection employs the locality term defined in Equation 10, which prioritizes assigning similar tasks to servers that have recently processed comparable workloads. This locality-aware assignment strategy effectively reduces the frequency of costly model switching operations and leverages warm-up effects from cached model weights.
6.3.3 Load Balance
Figure 10 demonstrates the superior load balancing performance of TORTA across all tested network topologies. The cumulative distribution functions reveal that TORTA consistently achieves the lowest load balance coefficients, with mean values of 0.765, 0.743, 0.755, and 0.745 for Abilene, Cost2, Gabriel, and Polska topologies respectively. The load balance coefficient $LB$ is defined as:
@@ 6.3.3 Load Balance
6.3.4 Details about Performance Metrics
@@
Figure 11 provides a detailed breakdown of response time components, revealing that TORTA achieves superior performance in both waiting times and inference times across all network topologies. The waiting times for TORTA range from 0.3 to 1.1 seconds across different topologies, representing significant reductions compared to baseline methods. This improvement in queuing delays can be attributed to TORTA’s proactive resource management and temporal-aware allocation strategy, which prevents task accumulation by anticipating future demand and pre-positioning resources accordingly.
@@
This threshold (0.5) suggests that the prediction module does not need to achieve exceptional accuracy to provide meaningful benefits. The 0.5 accuracy level corresponds to reasonable demand forecasting capabilities that can be obtained using standard time series prediction techniques, making TORTA practical for real-world deployment scenarios.
Even when prediction accuracy falls below the 0.5 threshold, TORTA’s performance degradation is gradual rather than catastrophic, indicating the robustness of the two-stage allocation scheme. The micro-level server selection component effectively compensates for macro-level prediction inaccuracies through real-time adaptation and intelligent task-server matching. This resilience ensures that TORTA remains competitive even when demand forecasting is challenging, while providing substantial performance gains when reasonable prediction accuracy is available.
7 Related Work
@@
8 Conclusion
This paper presents TORTA, a novel framework for distributed GPU inference optimization that addresses the fundamental limitations of existing reactive scheduling approaches. By integrating reinforcement learning with optimal transport theory and incorporating temporal dependencies into the allocation process, TORTA achieves significant improvements in response times, load balancing, and operational costs across diverse network topologies. The two-layer hierarchical architecture effectively decomposes the complex spatiotemporal optimization problem, enabling scalable solutions that consider both geographical resource distribution and temporal dynamics.
Experimental results demonstrate that TORTA consistently outperforms baseline methods across multiple metrics, achieving response time reductions of up to 15% and substantial improvements in load balancing coefficients. The framework exhibits robustness to prediction inaccuracies and maintains competitive performance even when demand forecasting accuracy is modest. As large language model deployment continues to scale globally, TORTA provides a practical and theoretically grounded approach to managing distributed GPU resources efficiently. Future work will focus on extending the framework to handle multi-tenant scenarios and investigating adaptive prediction mechanisms that can improve forecasting accuracy over time.
Appendix A Appendix: Theoretical Analysis of TARTO
@@
A.1 Upper Bound for Single-Timeslot Methods
@@
[…]
@@
For power cost optimality, since OT minimizes $\sum_{i,j}C_{i,j}P_{i,j}$ where $C_{i,j}$ incorporates regional electricity price differences, it naturally achieves minimum power cost. ∎
A.2 Switching Cost Convergence Analysis
@@
Proof.
Let $\Delta_{t}^{\mathcal{M}}=||A_{t}^{\mathcal{M}}-A_{t-1}^{\mathcal{M}}||_{F}^{2}$ denote the switching cost. We have:
$\mathbb{E}[\Delta_{t}^{\mathcal{M}}]=\mathbb{E}[||A_{t}^{\mathcal{M}}-A_{t-1}^{\mathcal{M}}||_{F}^{2}]$ (15)
$=\mathbb{E}[||A_{t}^{\mathcal{M}}||_{F}^{2}]+\mathbb{E}[||A_{t-1}^{\mathcal{M}}||_{F}^{2}]-2\mathbb{E}[\langle A_{t}^{\mathcal{M}},A_{t-1}^{\mathcal{M}}\rangle]$ (16)
@@
$\mathbb{E}[\langle A_{t}^{\mathcal{M}},A_{t-1}^{\mathcal{M}}\rangle]=\langle\mathbb{E}[A_{t}^{\mathcal{M}}],\mathbb{E}[A_{t-1}^{\mathcal{M}}]\rangle=||\mathbb{E}[A^{\mathcal{M}}]||_{F}^{2}$
@@
$\mathbb{E}[\Delta_{t}^{\mathcal{M}}]$ $=2\mathbb{E}[||A^{\mathcal{M}}||_{F}^{2}]-2||\mathbb{E}[A^{\mathcal{M}}]||_{F}^{2}$ (17)
@@
The key insight is that for methods satisfying the same constraints (demand and capacity), the variance is primarily determined by input distribution randomness rather than the optimization algorithm itself, leading to method-independent convergence. ∎
@@
A.3 TORTA’s Performance Guarantee
@@
$\mathcal{L}_{\epsilon}(s,\epsilon)$ $=\max\left(0,\frac{||B_{t}||_{F}-\epsilon_{\max}(s)}{\epsilon_{0}}\right)$ (19)
@@
    Single-timeslot performance control: $\mathbb{E}[||B_{t}||_{F}]\leq\epsilon$
@@
$\mathbb{E}[\Delta_{t}^{\text{RL}}]=\mathbb{E}[||A_{t}^{\text{RL}}-A_{t-1}^{\text{RL}}||_{F}^{2}]\leq\frac{K_{0}}{s}$
@@
$\delta R_{t}$ $\leq L_{R}||B_{t}||_{F}$ (21)
$\delta P_{t}$ $\leq L_{P}||B_{t}||_{F}$ (22)
where $L_{R},L_{P}$ are Lipschitz constants. The constraint $\mathbb{E}[||B_{t}||_{F}]\\ \leq\epsilon$ is enforced by $\mathcal{L}_{\epsilon}$.
@@
Therefore, $\text{Cost}_{\text{TORTA}}<\text{Cost}_{\text{Upper}}$, proving TORTA’s strict superiority. ∎

### v2

@@ V-B2 Reinforcement Learning with OT Supervision
$r_{t}=r_{OT}(A_{t},P^{*}_{t})+\lambda_{1}\cdot r_{smooth}(A_{t},A_{t-1})+\lambda_{2}\cdot r_{cost}(Q_{t})$ (3)
@@
The smoothness term $r_{smooth}(A_{t},A_{t-1})=-\|A_{t}-A_{t-1}\|_{F}^{2}$ penalizes large changes between consecutive allocation matrices, promoting temporal continuity and preventing abrupt load redistributions that can cause system instability. The cost term $r_{\text{cost}}(Q_{t})=-\frac{\|Q_{t}\|_{1}}{Q_{\text{max}}}$ utilizes the aggregate backlog across regional servers as a real-time proxy for system responsiveness, where $Q_{t}$ represents the vector of pending task counts and $Q_{\text{max}}$ denotes the maximum tolerable queue capacity. The weights $\lambda_{1}$ and $\lambda_{2}$ are set to prioritize OT alignment while ensuring sufficient temporal smoothness and load balance, with the relative weighting empirically tuned to achieve stable convergence during training.
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{PPO}}+\gamma\cdot\mathcal{L}_{\epsilon}(s,\epsilon)+\delta\cdot\mathcal{L}_{s}(s,\epsilon)$ (5)
@@
This design ensures that TORTA operates within the theoretical trade-off region where reduced switching costs outweigh single-timeslot optimality losses. The constraint parameters $\epsilon$ and $s$ must satisfy a specific relationship to guarantee theoretical performance advantages over any reactive method: $\frac{1-1/s}{\epsilon}>\frac{L_{R}+\beta L_{P}}{K_{0}}$, where $L_{R}$, $L_{P}$ are Lipschitz constants of response time and power cost functions, and $K_{0}$ is the baseline switching cost of reactive methods. The rigorous mathematical derivation of this relationship, detailed training procedures, and the formal proof are provided in Appendix.
V-C Micro-Level Allocation
@@
V-C1 Dynamic Server Activation Strategy
@@
V-C2 Greedy Task-Server Matching
@@
where $Recent_{tasks}^{(s)}$ represents tasks recently processed by server $s$, and $Similarity(\text{task}_{i},\text{task}_{j})=w_{m}\cdot\mathbb{I}[\text{model}_{i}=\text{model}_{j}]+w_{c}\cdot\cos(\text{embed}_{i},\text{embed}_{j})$ measures task similarity based on model type matching and input embedding cosine similarity, with weights $w_{m},w_{c}$.
@@
// Phase 1: Macro-Level Allocation
// Demand and Supply Normalization
@@ 3 $Prob_{i\rightarrow j}\leftarrow$ Row-normalize $P^{*}$;
// Reinforcement Learning Optimization
[…]
@@ 5 $A_{t}\leftarrow\pi_{\theta}(s_{t},F_{t},A_{t-1})$;
// Regional Task Distribution
6 for each task $i\in\mathcal{N}_{t}$ do
7  $region_{i}\leftarrow$ Sample from $A_{t}[origin(i),:]$;
8  Add task $i$ to regional queue $\mathcal{T}_{region_{i}}$;
@@
// Phase 2: Micro-Level Server Selection
@@
  // Dynamic Server Activation
11  $N_{target}^{(r,t)}\leftarrow$ Compute target active servers;
12  Update active server set $\mathcal{S}_{r}^{active}$ based on $N_{target}^{(r,t)}$;
  // Task-Server Matching
13  $\mathcal{T}_{r}\leftarrow$ Sort regional tasks by urgency (deadline proximity);
14
@@ 13  $\mathcal{T}_{r}\leftarrow$ Sort regional tasks by urgency (deadline proximity);
16     for each active server $s\in\mathcal{S}_{r}^{active}$ do
17        $Score(i,s)\leftarrow w_{1}\cdot Comp_{hw}(i,s)+w_{2}\cdot Comp_{load}(i,s)+w_{3}\cdot Comp_{locality}(i,s)$;
18
@@ 19    $s^{*}\leftarrow\arg\max_{s}Score(i,s)$;
20     $X_{i,r,s^{*}}^{t}\leftarrow 1$;
21     $Util_{s^{*}}\leftarrow Util_{s^{*}}+ResourceReq_{i}$;
22     $Queue_{s^{*}}\leftarrow Queue_{s^{*}}+1$;
23
24
// System State Update
25 Update utilization matrix $U_{t+1}$, queue lengths $Q_{t+1}$;
$A_{t-1}\leftarrow A_{t}$ return $X_{t}$, $s_{t+1}$
@@ VI-A Experimental Setup
Network Topology and Infrastructure. We evaluate our scheme across four distinct network topologies that represent different scales and geographical distributions of GPU infrastructure. The topologies range from small-scale regional deployments to large-scale global distributions (Table I.a), enabling us to assess scalability and performance across various infrastructure configurations.
Each server cluster contains heterogeneous GPU resources with different capabilities and capacities (see Table I.b). The infrastructure setup is designed to reflect real-world deployment scenarios where different GPU types are optimized for specific workload characteristics.
@@ VI-B Evaluation Metrics
TABLE I: Experimental Infrastructure Configuration
@@
Topo. Node B/W. (Gbps) Lat. (ms)
Abilene 12 10 25
Polska 12 10 45
Gabriel 25 15 80
Cost2 32 20 150
@@
VI-C Experimental Analysis
VI-C1 Response Time Performance Analysis
@@
Figure 9: Power cost and operational overhead comparison across different network topologies. Power costs are measured in thousands of dollars (\$K), while operational overhead represent the relative overhead (normalized units).
The response time distributions shown in Figure 8 demonstrate the effectiveness of our TORTA scheme across four different network topologies. TORTA consistently achieves the fastest average task completion times, with mean response times of 16.39s, 19.31s, 17.58s, and 19.19s for Abilene, Cost2, Gabriel, and Polska topologies respectively. This represents substantial improvements of 12.4% (vs. SkyLB: 18.72s), 10.5% (vs. SkyLB: 21.58s), 12.4% (vs. SkyLB: 20.07s), and 6.5% (vs. SkyLB: 20.53s) compared to the best-performing baseline method across each topology.
@@
VI-C2 Cost Optimization and Economic Efficiency
Figure 9 presents a comprehensive cost analysis comparing TORTA with baseline methods across four network topologies. TORTA consistently achieves the lowest power costs across all topologies: 12.5K (vs. SkyLB’s 14.3K, 12.6% reduction), 11.1K (vs. SkyLB’s 13.2K, 15.9% reduction), 10.7K (vs. SkyLB’s 12.8K, 16.4% reduction), and 14.1K (vs. SkyLB’s 15.2K, 7.2% reduction) for Abilene, Cost2, Gabriel, and Polska topologies respectively. Compared to other baselines, TORTA achieves similar improvements of 12-16% over SDIB and RR methods.
The optimal transport component enables TORTA to leverage geographical cost variations by routing tasks to regions with lower electricity prices while maintaining network latency constraints. This cost-aware allocation strategy is particularly effective in Cost2 and Gabriel topologies, where TORTA achieves the most substantial power cost reductions of 15.9% and 16.4% respectively.
For operational overhead, TORTA demonstrates dramatic improvements: 0.8 vs. 2.9 (72% reduction), 2.7 vs. 4.4 (39% reduction), 1.3 vs. 3.3 (61% reduction), and 2.3 vs. 3.4 (32% reduction) compared to SkyLB across the four topologies. Compared to other baselines, TORTA achieves 71-79% reductions vs. SDIB and 73-78% reductions vs. RR methods. The operational overhead metric represents normalized system transition costs (measured in planning units) that encompasses actual model loading, memory management, and server state changes, extending beyond the theoretical switching cost used in our analysis (which measures allocation matrix differences as a proxy). TORTA’s micro-level server selection employs proactive preheating and locality-aware assignment, while SkyLB and SDIB use cache-aware strategies and RR lacks such optimizations, explaining the varied overhead levels across methods.
VI-C3 Load Balance
Figure 10 demonstrates the superior load balancing performance of TORTA across all tested network topologies. TORTA consistently achieves the highest load balance coefficients with mean values of 0.765, 0.743, 0.755, and 0.745 for Abilene, Cost2, Gabriel, and Polska topologies respectively. Compared to the best-performing baseline SkyLB, this represents improvements of 4.4% (0.765 vs 0.733), 4.1% (0.743 vs 0.714), 3.6% (0.755 vs 0.729), and 4.2% (0.745 vs 0.715) across the four topologies. The improvements over other baselines are more substantial: 4.1-6.9% vs SDIB and 6.8-11.5% vs RR methods. The load balance coefficient $LB$ is defined as:
@@ VI-C3 Load Balance
VI-C4 Details about Performance Metrics
@@
Figure 11 provides a detailed breakdown of response time components, showing that TORTA achieves superior performance in both waiting times and inference times across all network topologies. The waiting times for TORTA range from 0.3 to 1.1 seconds compared to 1.2-2.4 seconds for baseline methods, representing reductions of approximately 50-75%. This improvement stems from TORTA’s proactive resource management and temporal-aware allocation strategy that prevents task accumulation.
@@
This threshold (0.4) suggests that the prediction module does not need to achieve exceptional accuracy to provide meaningful benefits. The 0.4 accuracy level corresponds to reasonable demand forecasting capabilities that can be obtained using standard time series prediction techniques, making TORTA practical for real-world deployment scenarios.
Even when prediction accuracy falls below the 0.4 threshold, TORTA’s performance degradation is gradual rather than catastrophic, indicating the robustness of the two-stage allocation scheme. The micro-level server selection component effectively compensates for macro-level prediction inaccuracies through real-time adaptation and intelligent task-server matching. This resilience ensures that TORTA remains competitive even when demand forecasting is challenging, while providing substantial performance gains when reasonable prediction accuracy is available.
VII Related Work
@@
VIII Conclusion
This paper presents TORTA, a novel scheme for distributed GPU inference optimization that addresses the fundamental limitations of existing reactive scheduling approaches. By integrating reinforcement learning with optimal transport theory and incorporating temporal dependencies into the allocation process, TORTA achieves improvements in response times, load balancing, and operational overhead across diverse network topologies. The two-layer hierarchical architecture effectively decomposes the complex spatiotemporal optimization problem, enabling scalable solutions that consider both geographical resource distribution and temporal dynamics.
Experimental results demonstrate that TORTA consistently outperforms baseline methods across multiple metrics, achieving response time reductions of up to 15% and substantial improvements in load balancing coefficients. The scheme exhibits robustness to prediction inaccuracies and maintains competitive performance even when demand forecasting accuracy is modest. However, several limitations remain: TORTA’s performance gains depend on prediction quality with diminishing returns below 40% accuracy, and the two-layer optimization introduces computational overhead for extremely large-scale deployments. Future work will focus on extending the framework to multi-tenant scenarios, investigating adaptive prediction mechanisms for improved forecasting accuracy, and developing lightweight variants for resource-constrained environments.
Appendix A Theoretical Analysis of TARTO
@@
A-A Upper Bound for Single-Timeslot Methods
@@
$\sum_{j}\tilde{w}_{j}$ $=\sum_{j}w^{\prime}_{j}\quad\text{(workload conservation)}$ (13)
$\sum_{j}\tilde{P}_{i,j}$ $=\mu_{t}^{(i)}\quad\text{(demand constraint)}$ (14)
@@
[…]
A-B Switching Cost Convergence Analysis
@@
Proof 2.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
A-C TORTA’s Performance Guarantee
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]